OpenStudy (anonymous):

what is x √(x+8)=√(3x+8)?

OpenStudy (anonymous):

solution?

OpenStudy (anonymous):

\[x = 1/3(-8 + 73/\sqrt[3]{-512 + 3i \sqrt{14097}} + \sqrt[3]{-512 + 3i \sqrt{14097}) }\]

OpenStudy (anonymous):

X is obviously not 0 its ~~1.1156999888680778

OpenStudy (anonymous):

There is a solution in the complex domain.

jhonyy9 (jhonyy9):

xsqrt(x+8)=sqrt(3x+8) squared both sides x2(x+8)=3x+8 x3+8x2-3x=8

jhonyy9 (jhonyy9):

- so do you have thought it the same ?